OpenStudy (anonymous):

Solve the system of equations using the substitution method 12x-4y=10 -3x-y=6

OpenStudy (anonymous):

the wolfram alpha thing didn't work on this one:( @misty1212

OpenStudy (misty1212):

put a comma between them

OpenStudy (texaschic101):

look at your 2nd equation....now solve for y

OpenStudy (anonymous):

i did

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=12x-4y%3D10%2C+-3x-y%3D6 worked for me

OpenStudy (anonymous):

ok thanks! I dunno what I did differently

OpenStudy (texaschic101):

-3x - y = 6 -3x - 6 = y now sub -3x - 6 in for y in the other equation. Then solve for x. Once x is found, sub that back into either equation to find y

OpenStudy (anonymous):

I typed that into my homework and it said it was incorrect :(

OpenStudy (anonymous):

so -3x - (-3x-6) =6

OpenStudy (anonymous):

oh now iget it thereis no solution

OpenStudy (texaschic101):

ok...let me show you we know that y = -3x - 6 now sub that into your 1st equation 12x - 4y = 10 12x - 4(-3x - 6) = 10 can you finish ?

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

the homework things said correct

OpenStudy (anonymous):

(-7/12, - 17/4)

OpenStudy (anonymous):

so yup no solution, thanks!

OpenStudy (texaschic101):

there is a solution

OpenStudy (anonymous):

Yw.

OpenStudy (texaschic101):

@allibricker12 gave the correct solution....(-7/12, - 17/4)